# Quantum Meets SAR: A Novel Range-Doppler Algorithm for Next-Gen Earth Observation

### v3

@@ Abstract
Synthetic Aperture Radar (SAR) plays a vital role in remote sensing due to its ability to capture high-resolution images regardless of weather conditions or daylight. However, to transform the raw SAR signals into interpretable imagery, advanced data processing techniques are essential. A widely used technique for this purpose is the Range Doppler Algorithm (RDA), which takes advantage of Fast Fourier Transform (FFT) to convert signals into the frequency domain for further processing. However, the computational cost of this approach becomes significant when dealing with large datasets. This paper presents a Quantum Range Doppler Algorithm (QRDA) that utilizes the Quantum Fourier Transform (QFT) to accelerate processing compared to the classical FFT. Furthermore, it introduces a quantum implementation of the Range Cell Migration Correction (RCMC) in the Fourier domain, a critical step in the RDA pipeline that realigns the received echoes so that the energy from a target is concentrated in a single range bin across all azimuth positions. The performance of the quantum RCMC is evaluated and compared against its classical counterpart, demonstrating the potential of quantum computing in advanced SAR imaging.
Index Terms:
@@
Quantum computing, Synthetic aperture radar (SAR), Range Doppler Algorithm (RDA), Quantum Fourier Transform (QFT), Range Cell Migration Correction (RCMC), Remote Sensing
footnotetext: These authors contributed equally to this work
I Introduction
@@
Synthetic Aperture Radar (SAR) is a powerful active imaging modality widely used in remote sensing due to its ability to operate independently of weather conditions and ambient lighting. In a typical SAR system, an airborne or spaceborne platform emits microwave pulses toward a target area and records the backscattered echoes. The raw data collected in this way encode information about the reflectivity properties of the scene, and transforming it into a usable image requires sophisticated processing to account for the motion of the radar and the geometry of wave propagation. The Range Doppler Algorithm (RDA) is one of the most widely used methods for the formation of SAR images. Its computational core relies on the Fast Fourier Transform (FFT) to convert the data into the frequency domain, where key operations such as range compression, Range Cell Migration Correction (RCMC), and azimuth compression are performed. However, the complexity of FFT $O(n\log n)$, although efficient for moderate datasets, becomes a bottleneck for large-scale or high-resolution SAR applications, where near-real-time processing is often desirable. In this work, we explore a quantum computing approach to accelerate the RDA by replacing classical FFT-based steps with their quantum counterparts. The Quantum Fourier Transform (QFT) offers a theoretical exponential speedup over the FFT, but this advantage is contingent on maintaining quantum coherence throughout the entire processing pipeline. Measuring quantum states halfway through (e.g., after QFT but before RCMC) collapses the quantum states, negating the speedup by losing all superposition information from before the measurement and limiting it to a singular result. Thus, a practical quantum RDA must integrate all critical steps—including RCMC—within the quantum domain to avoid decoherence and maximize computational gains[1]. To address these challenges, the contribution of this paper is two-fold:
@@
    A quantum-domain Range Cell Migration Correction (RCMC) gate, which is a crucial part in the Range Doppler algorithm
@@
II Fundamentals of Quantum Computing
II-A Classical vs. Quantum Information
Classical computation is based on bits as its fundamental unit of information, where each bit represents a binary state (0 or 1). For a system with $m$ distinct states, the minimum number of bits required is $n=[\log_{2}m]$. In contrast, quantum computation takes advantage of the qubit (quantum bit), which generalizes the classical bit by leveraging superposition and entanglement[2].
II-B The Qubit and Superposition
@@
where $|\alpha|^{2}$ and $|\beta|^{2}$ tell us the probability of finding |$\ket{\psi}$ in the states |$\ket{0}$ and |$\ket{1}$. When a qubit is measured, it will only be found to be in the state |$\ket{0}$ or the state |$\ket{1}$. This quality of superposition is at the core of quantum computing, allowing us to tap into more possibilities with a single qubit, along with giving us exponential scaling, since $n$ qubits can represent $2^{n}$ states simultaneously, compared to classical bits, which can only store one state at a time[2].
II-C Quantum States and Basis Representations
@@
with $\sum_{i}|c_{i}|^{2}=1$. The squared modulus of a given coefficient $c_{i}$ gives the probability that measurement finds the system in the state |$\ket{v_{i}}$[2].
II-D Quantum Operators and Unitarity
Quantum operators are linear transformations that act on states. An operator $\hat{A}$ maps |$\ket{\psi}$ to another state $\hat{A}\ket{\psi}=\ket{\phi}$. For quantum computation, the operators must be unitary, satisfying $U^{\dagger}U=I$, ensuring reversibility and probability conservation[2].
II-E Key Quantum Gates
Quantum gates manipulate qubit states analogously to classical logic gates but with additional capabilities (e.g., phase shifts, superposition)[2]. Two critical single-qubit gates are:
  • Hadamard gate (H) which creates superposition from computational basis states as shown in Equation 3:
    $H=\frac{1}{\sqrt{2}}\left(\begin{matrix}1&1\\ 1&-1\\ \end{matrix}\right)\text{, where }H\ket{0}=\ket{+},H\ket{1}=\ket{-}$ (3)
@@ II-E Key Quantum Gates
    $P=\left(\begin{matrix}1&0\\ 0&e^{i\theta}\\ \end{matrix}\right)\text{, with }P\ket{0}=\ket{0},P\ket{1}=e^{i\theta}\ket{1}$ (4)
III Range Doppler Algorithm
@@
III-A Classical RDA
The classical RDA consists of four stages[3] [4] as depicted in Figure 2:
@@ III-A Classical RDA
    Range Compression: The raw signal undergoes range compression through frequency-domain matched filtering as expressed in Equation 5:
@@
    where $G(f_{\tau})$ is the range reference function. This operation collapses all targets with identical slant ranges into single trajectories while preserving phase information.
@@
    Range Cell Migration Correction (RCMC): A phase correction compensates for slant-range variations caused by platform motion as shown in Equation 7
@@
    where $D(f_{\eta},V)=\sqrt{1-(\lambda f_{\eta}/2V)^{2}}$ is the Doppler compression factor, $V$ is the platform velocity, and $R_{0}$ is the reference range. This step aligns the target responses from curved trajectories to straight lines.
@@
    Azimuth Compression: Final focusing is achieved through azimuth matched filtering and transformation back to the time domain as shown in Equation 8:
@@
    The matched filter is given by: $H(f_{a})=e^{-j\phi(f_{a})}$, where $\phi(f_{a})$ is the phase history of the target in the azimuth frequency domain.
Figure 2: Range Doppler Algorithm: block diagram.
III-B Quantum RDA
Building upon the classical RDA framework, the proposed quantum implementation replaces FFT operations with the Quantum Fourier Transform (QFT), offering an exponential speedup in theory. This also means that the whole algorithm from encoding to measurement must be performed in the quantum domain to actually achieve this speedup[1]. The proposed approach is depicted in Figure 3 where the classical information is first encoded into the quantum domain using amplitude encoding. The next steps involve applying the range compression filter, the RCMC filter, and the azimuth compression filter after moving to the frequency domain and back using QFT (and inverse QFT) applied to the range and azimuth lines, respectively. More details are provided in the following.
Figure 3: Quantum Range Doppler Algorithm: proposed quantum circuit approach.
III-C Amplitude Encoding
@@
where |$\ket{i}$ are the computational basis states of an n-qubit system (for $N=2^{n}$). The coefficients $x_{i}$ are the amplitudes that encode the classical data. The vector must be normalized, that is, $\sum_{i}|x_{i}|^{2}=1$[5] [6].
IV QFT and QRCMC
IV-A Quantum Fourier Transform
@@
where $k$ represents the frequency components for an $n$-qubit system[7].
@@
The structure of the circuit reflects a recursive decomposition of the Fourier transform, with each qubit undergoing a Hadamard gate followed by progressively finer phase rotations conditioned on higher-order qubits, as illustrated in Figure 4[8].
@@ IV-C Circuit Realization
V Results
We verified our quantum RCMC implementation using the Qiskit state vector simulator (AerSimulator), which emulates an ideal noise-free quantum computer [9]. We compared the results with classical processing on Sentinel-1 SAR data (64×64 subset) [10]. As shown in Figure 5, the phase difference plot demonstrates near-perfect alignment between the quantum and classical outputs, confirming that the quantum implementation reproduces the expected classical behavior. The minimal discrepancies observed are attributable to numerical precision rather than algorithmic error. This test confirms the correctness of the proposed quantum approach.
Figure 5: The Phase Value Difference of the RDA Classical and Quantum approaches.
VI Conclusion

### v4

@@ Abstract
Synthetic Aperture Radar (SAR) plays a vital role in remote sensing due to its ability to capture high-resolution images regardless of weather conditions or daylight. However, to transform the raw SAR signals into interpretable imagery, advanced data processing techniques are essential. A widely used technique for this purpose is the Range Doppler Algorithm (RDA), which takes advantage of Fast Fourier Transform (FFT) to convert signals into the frequency domain for further processing. However, the computational cost of this approach becomes significant when dealing with large datasets. This paper presents a Quantum Range Doppler Algorithm (QRDA) that utilizes the Quantum Fourier Transform (QFT) to accelerate processing compared to the classical FFT. Furthermore, it introduces a quantum implementation of the Range Cell Migration Correction (RCMC) in the Fourier domain, a critical step in the RDA pipeline that realigns the received echoes so that the energy from a target is concentrated in a single range bin across all azimuth positions. The performance of the quantum RCMC is evaluated and compared against its classical counterpart, as an isolated operation, and then as part of the full classical pipeline, demonstrating the potential of quantum computing in advanced SAR imaging.
keywords:
@@
Quantum computing, Synthetic aperture radar (SAR), Range Doppler Algorithm (RDA), Quantum Fourier Transform (QFT), Range Cell Migration Correction (RCMC), Remote Sensing.
1 Introduction
@@
Synthetic Aperture Radar (SAR) is a powerful active imaging modality widely used in remote sensing due to its ability to operate independently of weather conditions and ambient lighting. In a typical SAR system, an airborne or spaceborne platform emits microwave pulses toward a target area and records the backscattered echoes. The raw data collected in this way encode information about the reflectivity properties of the scene, and transforming it into a usable image requires sophisticated processing to account for the motion of the radar and the geometry of wave propagation. The Range Doppler Algorithm (RDA) is one of the most widely used methods for the formation of SAR images. Its computational core relies on the Fast Fourier Transform (FFT) to convert the data into the frequency domain, where key operations such as range compression, Range Cell Migration Correction (RCMC), and azimuth compression are performed. However, the complexity of FFT $O(n\log n)$, although efficient for moderate datasets, becomes a bottleneck for large-scale or high-resolution SAR applications, where near-real-time processing is often desirable. In this work, we explore a quantum computing approach to accelerate the RDA by replacing classical FFT-based steps with their quantum counterparts. The Quantum Fourier Transform (QFT) offers a theoretical exponential speedup over the FFT, but this advantage is contingent on maintaining quantum coherence throughout the entire processing pipeline. Measuring quantum states halfway through (e.g., after QFT but before RCMC) collapses the quantum states, negating the speedup by losing all superposition information from before the measurement and limiting it to a singular result. Thus, a practical quantum RDA must integrate all critical steps, including RCMC, within the quantum domain to avoid decoherence and maximize computational gains [Agency, 2023].
To address these challenges, the contribution of this paper is three-fold as follows:
@@
    A quantum-domain Range Cell Migration Correction (RCMC) gate, which is a crucial part in the Range Doppler algorithm. Figure 1 shows the proposed RCMC quantum circuit.
  3. 3.
    Evaluation of the proposed RCMC gate on real SAR data in two setups: (i) isolation on a 64×64 pixel subset Sentinel-1 SAR data and (ii) integration in a full classical RDA.
@@
The rest of this paper is organized as follows: Section 2 introduces the basic principles of quantum computing, including qubits, gates, and unitary operators. Section 3 reviews the classical Range Doppler Algorithm (RDA). Section 4 presents the proposed quantum circuit implementation of the RDA, with emphasis on the Quantum Fourier Transform (QFT) and the Quantum Range Cell Migration Correction (QRCMC) gate. Section 5 evaluates the performance and correctness of our quantum approach through simulations on real SAR data in isolation and in integration in a full classical RDA. Finally, Section 6 summarizes our findings and outlines directions for future research in quantum SAR processing.
2 Fundamentals of Quantum Computing
Classical computation is based on bits as its fundamental unit of information, where each bit represents a binary state (0 or 1). For a system with $m$ distinct states, the minimum number of bits required is $n=[\log_{2}m]$. In contrast, quantum computation takes advantage of the qubit (quantum bit), which generalizes the classical bit by leveraging superposition and entanglement [McMahon, 2007].
2.1 The Qubit and Superposition
@@
where $|\alpha|^{2}$ and $|\beta|^{2}$ tell us the probability of finding |$\ket{\psi}$ in the states |$\ket{0}$ and |$\ket{1}$. When a qubit is measured, it will only be found to be in the state |$\ket{0}$ or the state |$\ket{1}$. This quality of superposition is at the core of quantum computing, allowing us to tap into more possibilities with a single qubit, along with giving us exponential scaling, since $n$ qubits can represent $2^{n}$ states simultaneously, compared to classical bits, which can only store one state at a time [McMahon, 2007].
2.2 Quantum States and Basis Representations
@@
with $\sum_{i}|c_{i}|^{2}=1$. The squared modulus of a given coefficient $c_{i}$ gives the probability that measurement finds the system in the state |$\ket{v_{i}}$ [McMahon, 2007].
2.3 Quantum Operators and Unitarity
Quantum operators are linear transformations that act on states. An operator $\hat{A}$ maps |$\ket{\psi}$ to another state $\hat{A}\ket{\psi}=\ket{\phi}$. For quantum computation, the operators must be unitary, satisfying $U^{\dagger}U=I$, ensuring:
  • Reversibility: Reversible operations preserve all quantum information throughout the computation, which is vital not only for algorithmic correctness but also for maintaining quantum coherence. Moreover, irreversible operations would imply loss of information, which contradicts the deterministic evolution of isolated quantum systems.
  • Probability conservation: Unitary operators preserve the inner product structure of Hilbert space, which ensures that the norm of the state vector remains constant during evolution [McMahon, 2007].
2.4 Key Quantum Gates
Quantum gates manipulate qubit states analogously to classical logic gates but with additional capabilities (e.g., phase shifts, superposition) [McMahon, 2007]. Two critical single-qubit gates are: Hadamard gate (H) which creates superposition from computational basis states as shown in Equation 3:
$H=\frac{1}{\sqrt{2}}\left(\begin{matrix}1&1\\ 1&-1\\ \end{matrix}\right)$ (3)
where $H\ket{0}=\ket{+}$
$|H\ket{1}=\ket{-}$
@@ 2.4 Key Quantum Gates
$P=\left(\begin{matrix}1&0\\ 0&e^{i\theta}\\ \end{matrix}\right)$ (4)
where $P\ket{0}=\ket{0}$
$P\ket{1}=e^{i\theta}\ket{1}$
3 Range Doppler Algorithm
@@
3.1 Classical RDA
The classical RDA consists of four stages [Dastgir, 2007] [Agency, 2019] as depicted in Figure 2:
@@ 3.1 Classical RDA
    Range Compression: is the first major processing step in the Range-Doppler Algorithm (RDA), and its primary goal is to enhance range resolution by concentrating the dispersed pulse energy reflected from each target into a narrow peak. This step significantly improves the signal-to-noise ratio (SNR) which quantifies how much a signal stands out from background noise, and prepares the data for further processing such as Range Cell Migration Correction (RCMC) and azimuth compression.
    In most practical SAR systems, this operation is performed via frequency-domain matched filtering. Instead of correlating the received signal with the time, reversed transmitted pulse directly in the time domain, a computationally intensive process—this filtering is efficiently implemented in the frequency domain using the convolution theorem. The signal is first transformed into the frequency domain using the FFT, then multiplied by a range reference function $G(f_{\tau})$, and finally transformed back using the inverse FFT:
@@
    where $G(f_{\tau})$ is the range reference function, $(\tau,\eta)$ is the received raw signal, $\tau$ is the fast time (range), $\eta$ is the slow time (azimuth). This operation collapses all targets with identical slant ranges into single trajectories while preserving phase information.
    The matched filter maximizes the SNR by aligning the received signal with the known transmitted pulse in phase, effectively compressing the wide-bandwidth chirp signal into a sharp spike in time. As a result, echoes from targets at the same slant range but different azimuth positions become concentrated along single trajectories in the 2D signal space, while maintaining phase information crucial for later steps such as interferometry or Doppler processing.
    From a signal processing perspective, this filtering suppresses range ambiguities and background noise, making it easier to detect true target reflections. The output signal after range compression reveals the relative distances to the targets by the locations of the resulting peaks in the time domain. The peak amplitudes reflect the intensity of the backscattered energy, which is influenced by both the target’s radar cross-section and its distance from the radar.
    Visually, the range-compressed data appears as a superposition of hyperbolic arcs in the $(\tau,\eta)$ plane, where each arc corresponds to a ground target. These arcs represent the geometric locus of reflections from a target as the radar platform moves through the synthetic aperture.
    Figure 2: Block diagram of the classical Range Doppler Algorithm (RDA) used in Synthetic Aperture Radar (SAR) imaging.
@@
    Range Cell Migration Correction (RCMC): is a critical step in the Range Doppler Algorithm (RDA), addressing the range migration effect caused by the relative motion between the radar platform and the ground targets during the synthetic aperture. Due to this relative motion, echoes from a single point target do not remain in a fixed range bin over the aperture time but instead migrate across multiple bins. If uncorrected, this effect degrades image focus and resolution.
    RCMC realigns these dispersed echoes by applying a phase correction in the range frequency domain, effectively remapping the signal so that each target’s energy is concentrated in a single range cell. This correction allows subsequent azimuth compression to be applied correctly, ensuring that the resulting image is sharp and geometrically accurate.
    Mathematically, the correction is applied as a linear phase shift in the range frequency domain: 7
@@
    where $D(f_{\eta},V)=\sqrt{1-(\lambda f_{\eta}/2V)^{2}}$ is the Doppler compression factor, $V$ is the platform velocity, and $R_{0}$ is the reference range, $f_{r}$ is the range frequency, and $f_{\eta}$ is the azimuth (Doppler) frequency,. This exponential term serves as a frequency-dependent phase multiplier that corrects for the slant-range curvature introduced by platform motion. The Doppler compression factor $D(f_{\eta},V)$ accounts for the nonlinearity in the trajectory of target returns. As a result, RCMC flattens the curved migration path of target echoes, allowing them to be coherently integrated in the azimuth direction.
    In implementation, this correction is often performed in the frequency domain using FFTs, as it allows for efficient convolution. The phase shift translates into a range shift in the time domain, aligning all backscattered signals from a single target.
    An intuitive way to understand RCMC is to imagine the radar moving past a stationary object. At the start of the aperture, the object appears in one range bin, but as the radar moves, the geometry changes and the object’s echo appears to ”drift” through other range bins. RCMC ”undoes” this drift, essentially straightening a curved path into a linear one.
@@
    Azimuth Compression:is the final step in the Range-Doppler Algorithm (RDA), responsible for focusing the radar returns in the azimuth (along-track) direction and producing the final high-resolution SAR image. Just as range compression improves resolution in the range dimension by applying matched filtering, azimuth compression performs an analogous operation in the Doppler (frequency) domain of azimuth.
    After the Range Cell Migration Correction (RCMC) step, the radar data is organized in the range-Doppler domain. Each row of this two-dimensional dataset corresponds to a specific slant range, and each column represents a Doppler frequency (related to azimuth). The azimuth signal for each range bin must now be focused to concentrate the target energy into a sharp point in azimuth.
    This is achieved through matched filtering in the azimuth frequency domain. The process applies a complex exponential filter designed to reverse the phase evolution imparted by the relative motion between the radar and the target. The phase history $\phi(f_{a})$ of a stationary target during the synthetic aperture is typically quadratic, and correcting it allows for coherent integration of the received echoes. The operation is expressed mathematically as: 8:
@@
    $s_{2}(\tau,f_{\eta})$ is the signal after range compression and RCMC, $H(f_{\eta})=e^{-j\phi(f_{\eta})}$ is the azimuth matched filter, $\eta$ is azimuth (slow) time, $f_{\eta}$ is the azimuth (Doppler) frequency.
    The matched filter is given by: $H(f_{a})=e^{-j\phi(f_{a})}$, where $\phi(f_{a})$ is the phase history of the target in the azimuth frequency domain The matched filter $H(f_{\eta})$ cancels out the phase modulation imposed by the radar-target geometry, which is typically parabolic due to the curved relative motion of the radar platform. By applying this filter, the algorithm aligns the phases of all echo contributions from the same target, thereby concentrating their energy and enhancing image sharpness in azimuth.
    Following the application of the filter, an inverse FFT (IFFT) is used to transform the frequency-domain signal back to the azimuth time domain. This step completes the formation of the SAR image, delivering a focused two-dimensional representation of the scene with high resolution in both range and azimuth.
    The effectiveness of azimuth compression is crucial for fine resolution and image clarity. Without it, the energy of each target would be spread out along the azimuth dimension, resulting in blurred and unfocused images. This step not only ensures geometric accuracy but also plays a pivotal role in improving contrast and enabling applications such as change detection or interferometric SAR (InSAR).
3.2 Quantum RDA
Building upon the classical RDA framework, the proposed quantum implementation replaces FFT operations with the Quantum Fourier Transform (QFT), offering an exponential speedup in theory. This also means that the whole algorithm from encoding to measurement must be performed in the quantum domain to actually achieve this speedup [Agency, 2023]. The proposed approach is depicted in Figure 3 where the classical information is first encoded into the quantum domain using amplitude encoding. The next steps involve applying the range compression filter, the RCMC filter, and the azimuth compression filter after moving to the frequency domain and back using QFT (and inverse QFT) applied to the range and azimuth lines, respectively. More details are provided in the following.
Figure 3: Quantum Range-Doppler Algorithm (QRDA): Proposed quantum circuit implementation of the classical Range-Doppler Algorithm (RDA).
3.3 Amplitude Encoding
@@
where |$\ket{i}$ are the computational basis states of an n-qubit system (for $N=2^{n}$). The coefficients $x_{i}$ are the amplitudes that encode the classical data. The vector must be normalized, that is, $\sum_{i}|x_{i}|^{2}=1$ [Date, 2024] [Ranga, 2024].
4 QFT and QRCMC
After successfully encoding the SAR raw data into quantum states, we can proceed with the application of the Quantum Fourier Transform (QFT) and the Quantum Range Cell Migration Correction (QRCMC). The QFT allows efficient transformation into the frequency domain, while the QRCMC applies precise phase corrections to compensate for range cell migration due to target motion.
4.1 Quantum Fourier Transform
@@
where $k$ represents the frequency components for an $n$-qubit system [Qiskit, 2023].
@@
The structure of the circuit reflects a recursive decomposition of the Fourier transform, with each qubit undergoing a Hadamard gate followed by progressively finer phase rotations conditioned on higher-order qubits, as illustrated in Figure 4 [Jaffali, 2022].
@@ 4.3 Circuit Realization
Figure 5: Phase difference between classical and quantum RDA outputs for a 64×64 Sentinel-1 subset. The close match confirms the quantum implementation’s accuracy.
@@
5 Results
We validated our quantum RCMC (Range Cell Migration Correction) implementation using the Qiskit AerSimulator, a state vector simulator that emulates an ideal, noise-free quantum environment [Javadi-Abhari, 2024]. To assess correctness, we compared the quantum outputs against results obtained from classical processing on a 64×64 pixel subset of Sentinel-1 SAR data [Hall, 2025].
As shown in Figure 5, the phase difference plot reveals near-perfect alignment between the classical and quantum outputs. This confirms that our quantum algorithm reproduces the expected classical behavior. The minor discrepancies observed can be attributed to numerical precision limitations inherent in floating-point calculations, rather than errors in algorithmic logic. This test demonstrates the functional correctness of the proposed quantum RCMC algorithm in isolation.
To evaluate the quantum RCMC’s real-world applicability, we integrated the QRCMC gate into a full classical Range-Doppler Algorithm (RDA) processing chain. For this experiment, we selected a smaller 8×8 subset of real Sentinel-1 RAW data. This size was chosen deliberately to remain within the practical limits of current quantum simulators, which are constrained by exponential growth in memory and computation time as the number of qubits increases. Despite the reduced dimensionality, this test case is sufficient to validate the algorithm’s end-to-end correctness.
As illustrated in Figure 6, the resulting phase difference again demonstrates excellent alignment between the hybrid (quantum-integrated) and purely classical outputs. Similar to the isolated test case, any observed discrepancies are minor and attributable to floating-point numerical precision, not algorithmic faults.
These results confirm that the quantum RCMC component can be correctly embedded within a full RDA workflow and operate in conjunction with classical stages, thereby demonstrating its functional reliability under realistic conditions.
Figure 6: Phase difference between classical and quantum-integrated RDA outputs for an 8×8 Sentinel-1 subset. The high degree of alignment confirms successful integration of the QRCMC module.
6 Conclusion